-obtain wavelengths for the hydrogen emission lines and then calculate the energy levels according to bohr's theory

- determine wavelengths in nm associated with following transitions of the H atom:

nf = 1 with ni = 2, 3, 4

nf = 2 with ni = 3, 4, 5

nf = 3 with ni = 4, 5, 6

Hydrogen יויויין יויויויויויויוי 400 500 600 700 Fig e: Hydrogen emission spectra for calculations

1 1 Rydberg Equation: -R, a n 1 1 91 nm For hydrogen atom: 2 n 2 ni λ
Answer #1

to Bohrs theory According I = RH ( 11 where RA CRydberg constant) = 10973731 ma! So for ng =1, with mi=2 1 nm = 10973731 (1-PO Paschen Series with For nf ni = 4 w - RH lå - to D = RH x 0.0486 9. I X10 1875.03 mm. (Am) 10973731 x 0.0486 nm nf=3; ni=For, hert Cart - 1/2) 13.6 ev -1511er cam) heRH Eu R+ (a) - 13.6 16 0.85 ev (am] ev ev Es 13.6 25 hert ca 1 / 2 11 -0.544 ev.